# Modeling Cell Dynamics and Interactions with Unbalanced Mean Field Schrödinger Bridge

### v2

@@ 5.1 Simulating ODEs: Random Batch Methods
In ODE simulation, the computation complexity is $\mathcal{O}(N^{2})$ due to the cellular interaction term. To speed up simulation, we adopt the Random Batch Methods (RBMs) (Jin et al., 2020, 2021; Jin & Li, 2022), which transforms the interaction among all particles into interaction with particles within random grouping only. The algorithm reduces the computational complexity to $\mathcal{O}((p-1)N)$, where $p$ is the number in the batch $\mathcal{C}_{p}$. Assuming $\mathbf{v}(\mathbf{x},t)$ and $\Phi$ satisfy certain conditions, it has the convergence result such that $\mathcal{W}_{2}\left(\tilde{\mu}_{N}^{(1)}(t),\mu_{N}^{(1)}(t)\right)\leq C\sqrt{\tau}$ (Jin et al., 2020, 2021), where $\tau$ is the step size. The $\tilde{\mu}_{N}^{(1)}(t)$ is the empirical distribution produced by Algorithm 1 and $\mu_{N}^{(1)}(t)$ is the distribution by simulating the original ODEs.
$m\in 1:[T/\tau]$
[…]
@@ 5.2 Reformulating the Loss with Weighted Particle Simulation
The total loss to solve Theorem 4.1 is composed of three parts, e.g., the energy loss, the reconstruction loss, and the Fokker-Planck constraint such that $\mathcal{L}=\mathcal{L}_{\text{Energy}}+\lambda_{r}\mathcal{L}_{\text{Recons}}+\lambda_{f}\mathcal{L}_{\text{FP}}.$ Here $\mathcal{L}_{\text{Energy}}$ loss promotes the least action principle of transition energy, $\mathcal{L}_{\text{Recons}}$ promotes the matching loss. i.e., $\rho(\mathbf{x},1)=\nu_{1}$, and $\mathcal{L}_{\text{FP}}$ promotes the three neural networks that satisfy the Fokker-Planck equation constraint. We reformulate the loss terms through weighted particle representation and an RBM-based Neural ODE solver.
@@
Generalizing the idea in (Sha et al., 2024), the energy loss in CytoBridge is equivalent to
$\mathbb{E}_{\mathbf{x_{0}}\sim\rho_{0}}\int_{0}^{T}\left[\frac{1}{2}\left\|{\mathbf{v}}_{\theta}(\mathbf{x}(t),t)\right\|_{2}^{2}+\frac{1}{2}\left\|\nabla_{\mathbf{x}}s_{\theta}\right\|_{2}^{2}+\langle{\mathbf{v}}_{\theta}(\mathbf{x}(t),t),s_{\theta}\rangle+\alpha\Psi\left(g_{\theta}\right)\right]w_{\theta}(t)\mathrm{d}t,$
where $w_{\theta}(t)=\exp\left({\int_{0}^{t}g_{\theta}(\mathbf{x}(t),s)\mathrm{d}s}\right)$ and $\mathbf{x}(t)$ satisfies $\mathrm{d}\mathbf{x}^{i}/\mathrm{d}t=\mathbf{v}(\mathbf{x}^{i},t)-\frac{1}{N-1}\sum_{j\neq i}{k(\mathbf{x}^{i},\mathbf{x}^{j})}w_{j}\nabla_{\mathbf{x}^{i}}\Phi(\mathbf{x}^{i}-\mathbf{x}^{j})$. However, direct optimization of this term is challenging due to the involvement of the inner product $\langle\mathbf{v}_{\theta}(\mathbf{x}(t),t),s_{\theta}\rangle$ which introduces mutual dependencies between the optimization of $\mathbf{v}$ and $s_{\theta}$. To address this issue and simplify the computation, we adopt an upper bound of the energy for training purposes (Section A.5).
@@
The reconstruction loss aims to align the final generated density to the true data density. Here we need to consider the unbalanced effect, so we use the unbalanced optimal transport to align it. $\mathcal{L}_{\text{Recons}}=\lambda_{m}\mathcal{L}_{\text{Mass}}+\lambda_{d}\mathcal{L}_{\text{OT}}.$ The $\mathcal{L}_{\text{Mass}}$ is used to obtain the cell weights and align the cell number/ mass in the datasets. We then use the weights to normalize the distribution and apply $\mathcal{L}_{\text{OT}}$ to match the distribution. In this work, we employ the local mass matching strategy from (Zhang et al., 2025a). Specifically, the trajectory mapping function $\phi_{\theta}^{\mathbf{v}}$ predicts particle coordinates $\widehat{A}_{1},\ldots,\widehat{A}_{T-1}$ from an initial set $A_{0}$ over time indices $\mathcal{T}$, governed by $\mathrm{d}\mathbf{x}/\mathrm{d}t=\mathbf{v}_{\theta}$ if no interaction is considered, or with a modified velocity $\widetilde{\mathbf{v}}$ incorporating the interaction potential $\Phi$: $\widetilde{\mathbf{v}}(\mathbf{x}^{i},t)={\mathbf{v}}(\mathbf{x}^{i},t)-\frac{1}{N-1}\sum_{j\neq i}{k(\mathbf{x}^{i},\mathbf{x}^{j})}\nabla_{\mathbf{x}^{i}}\Phi(\mathbf{x}^{i}-\mathbf{x}^{j})$. Additionally, a weight mapping function $\phi_{\theta}^{g}$ models particle weights via $\mathrm{d}\log w_{i}(t)/\mathrm{d}t=g_{\theta}(\mathbf{x}_{i}(t),t)$, starting from initial weights $w_{i}(0)=1/N$ where $N$ represents batch size. The mass matching loss $\mathcal{L}_{\text{Mass}}$ is composed of two terms. The first term is defined as the local mass matching loss $\mathcal{L}_{\text{Local Mass}}=\sum_{k=1}^{T-1}M_{k}$, where $M_{k}$ quantifies the error between predicted weights $w_{i}(t_{k})$ and target weights based on the cardinality of mapped points. This term provides fine-grained guidance on the growth network. Besides, the second term is defined as the global mass matching loss $\mathcal{L}_{\text{Global Mass}}=\sum_{k=1}^{T-1}G_{k}$, where $G_{k}$ is used to align the change of weights in total. An optimal transport loss $\mathcal{L}_{\text{OT}}=\sum_{k=1}^{T-1}\mathcal{W}_{2}(\hat{\mathbf{w}}^{k},\mathbf{w}(t_{k}))$ further aligns the predicted and observed distributions. Details can be found in Section A.4.
@@
CytoBridge aims to train four neural networks to model cell dynamics and interactions. To stabilize the training procedure, we leverage a two-phase training strategy. In the pre-training phase, we seek to provide a suitable initialization of these four neural networks. We first initialize $g_{\theta}$ and $\mathbf{v}_{\theta}$ without the interaction term to provide an approximated matching. Then we train ${\Phi}_{\theta}$ with fixed $g_{\theta}$, leading to refined dynamics. The score network $s_{\theta}$ is trained based on conditional flow matching. In the training phase, these initialized networks are further refined by minimizing the proposed total loss. We summarize the training procedure in Algorithm 2 and Appendix A. We conduct ablation studies on different components of our training procedure in Section B.6. The selection of loss weighting is discussed in Section C.2.
@@
Datasets $A_{0},\ldots,A_{T-1}$, batch size $N$, ODE iteration $n_{\text{ode}}$, log density iteration $n_{\text{log-density}}$rained neural ODE $\mathbf{v}_{\theta}$, growth function $g_{\theta}$, score network $s_{\theta}$and interaction ${\Phi}_{\theta}$.
1:
[…]
3:$\hat{A}_{t+1}\leftarrow\phi_{\theta}^{\mathbf{v}}(\hat{A}_{t},t+1)$, $w(\hat{A}_{t+1})\leftarrow\phi_{\theta}^{g}(w(\hat{A}_{t}),t+1)$
[…]
5:$\hat{A}_{t+1}\leftarrow\phi_{\theta}^{\widetilde{\mathbf{v}}}(\hat{A}_{t},t+1)$, $w(\hat{A}_{t+1})\leftarrow\phi_{\theta}^{g}(w(\hat{A}_{t}),t+1)$
[…]
7:$(\mathbf{x}_{0},\mathbf{x}_{1})\sim q(\mathbf{x}_{0},\mathbf{x}_{1})$, $t\sim\mathcal{U}(0,1)$, $\textbf{x}\sim p(\mathbf{x},t\mid\mathbf{x}_{0},\mathbf{x}_{1})$ using $\hat{A}_{0},\ldots,\hat{A}_{T-1}$
8:$\mathcal{L}_{\text{score}}\leftarrow\|\lambda_{s}\nabla_{\mathbf{x}}s_{\theta}(\mathbf{x},t)+\bm{\epsilon}_{1}\|_{2}^{2}$; update $s_{\theta}$ w.r.t. $\mathcal{L}_{\text{score}}$EndFor
9:
10:Training Phase:
[…]
12:$\hat{A}_{t+1}\leftarrow\phi_{\theta}^{\widetilde{\mathbf{v}}}(\hat{A}_{t},t+1)$, $w(\hat{A}_{t+1})\leftarrow\phi_{\theta}^{g}(w(\hat{A}_{t}),t+1)$
[…]
@@
15:$\mathcal{L}_{\text{Recons}}\leftarrow\lambda_{m}(M_{t}+G_{t})+\lambda_{d}\mathcal{W}_{2}(\hat{\mathbf{w}}^{t},\mathbf{w}(t))$
@@
16:$\mathcal{L}_{\text{FP}}\leftarrow\left\|\partial_{\tau}\rho_{\theta}(\mathbf{x},\tau)+\nabla_{\mathbf{x}}\cdot\left(\rho_{\theta}(\mathbf{x},\tau)\widetilde{\mathbf{v}}_{\theta}(\mathbf{x},\tau)\right)-g_{\theta}(\tau)\rho_{\theta}(\mathbf{x},\tau)\right\|+\lambda_{w}\left\|\rho_{\theta}(\mathbf{x},0)-p_{0}(\mathbf{x})\right\|$

### v3

@@ 5.1 Simulating ODEs: Random Batch Methods
In ODE simulation, the computation complexity is $\mathcal{O}(N^{2})$ due to the cellular interaction term. To speed up simulation, we adopt the Random Batch Methods (RBMs) (Jin et al., 2020, 2021; Jin and Li, 2022), which transforms the interaction among all particles into interaction with particles within random grouping only. The algorithm reduces the computational complexity to $\mathcal{O}((p-1)N)$, where $p$ is the number in the batch $\mathcal{C}_{p}$. Assuming $\mathbf{v}(\mathbf{x},t)$ and $\Phi$ satisfy certain conditions, it has the convergence result such that $\mathcal{W}_{2}\left(\tilde{\mu}_{N}^{(1)}(t),\mu_{N}^{(1)}(t)\right)\leq C\sqrt{\tau}$ (Jin et al., 2020, 2021), where $\tau$ is the step size. The $\tilde{\mu}_{N}^{(1)}(t)$ is the empirical distribution produced by Algorithm˜1 and $\mu_{N}^{(1)}(t)$ is the distribution by simulating the original ODEs.
1: for $m\in 1:[T/\tau]$ do
Randomly divide $\{1,2,\cdots,N=np\}$ into $n$ batches
2:   for each batch do
Update particles with $\mathrm{d}\mathbf{X}_{t}^{i}=\mathbf{v}(\mathbf{X}_{t}^{i},t)\mathrm{d}t-\frac{1}{p-1}\sum\limits_{j\neq i,j\in\mathcal{C}_{p}}^{N}{k_{i,j}w_{j}}\nabla_{\mathbf{x}}\Phi(\mathbf{X}_{t}^{i}-\mathbf{X}_{t}^{j})\mathrm{d}t$
@@ 5.2 Reformulating the Loss with Weighted Particle Simulation
The total loss to solve Theorem˜4.1 is composed of three parts, e.g., the energy loss, the reconstruction loss, and the Fokker-Planck constraint such that $\mathcal{L}=\mathcal{L}_{\text{Energy}}+\lambda_{r}\mathcal{L}_{\text{Recons}}+\lambda_{f}\mathcal{L}_{\text{FP}}.$ Here $\mathcal{L}_{\text{Energy}}$ loss promotes the least action principle of transition energy, $\mathcal{L}_{\text{Recons}}$ promotes the matching loss. i.e., $\rho(\mathbf{x},1)=\nu_{1}$, and $\mathcal{L}_{\text{FP}}$ promotes the three neural networks that satisfy the Fokker-Planck equation constraint. We reformulate the loss terms through weighted particle representation and an RBM-based Neural ODE solver.
@@
Generalizing the idea in (Sha et al., 2024), the energy loss in CytoBridge is equivalent to $\mathbb{E}_{\mathbf{x_{0}}\sim\rho_{0}}\int_{0}^{T}\left[\frac{1}{2}\left\|{\mathbf{v}}_{\theta}(\mathbf{x}(t),t)\right\|_{2}^{2}+\frac{1}{2}\left\|\nabla_{\mathbf{x}}s_{\theta}\right\|_{2}^{2}+\langle{\mathbf{v}}_{\theta}(\mathbf{x}(t),t),s_{\theta}\rangle+\alpha\Psi\left(g_{\theta}\right)\right]w_{\theta}(t)\mathrm{d}t,$ where $w_{\theta}(t)=\exp\left({\int_{0}^{t}g_{\theta}(\mathbf{x}(t),s)\mathrm{d}s}\right)$ and $\mathbf{x}(t)$ satisfies $\mathrm{d}\mathbf{x}^{i}/\mathrm{d}t=\mathbf{v}(\mathbf{x}^{i},t)-\frac{1}{N-1}\sum_{j\neq i}{k(\mathbf{x}^{i},\mathbf{x}^{j})}w_{j}\nabla_{\mathbf{x}^{i}}\Phi(\mathbf{x}^{i}-\mathbf{x}^{j})$. However, direct optimization of this term is challenging due to the involvement of the inner product $\langle\mathbf{v}_{\theta}(\mathbf{x}(t),t),s_{\theta}\rangle$ which introduces mutual dependencies between the optimization of $\mathbf{v}$ and $s_{\theta}$. To address this issue and simplify the computation, we adopt an upper bound of the energy for training purposes (Section˜A.5).
@@
The reconstruction loss aims to align the final generated density to the true data density. Here we need to consider the unbalanced effect, so we use the unbalanced optimal transport to align it. $\mathcal{L}_{\text{Recons}}=\lambda_{m}\mathcal{L}_{\text{Mass}}+\lambda_{d}\mathcal{L}_{\text{OT}}.$ The $\mathcal{L}_{\text{Mass}}$ is used to obtain the cell weights and align the cell number/ mass in the datasets. We then use the weights to normalize the distribution and apply $\mathcal{L}_{\text{OT}}$ to match the distribution. In this work, we employ the local mass matching strategy from (Zhang et al., 2025a). Specifically, the trajectory mapping function $\phi_{\theta}^{\mathbf{v}}$ predicts particle coordinates $\widehat{A}_{1},\ldots,\widehat{A}_{T-1}$ from an initial set $A_{0}$ over time indices $\mathcal{T}$, governed by $\mathrm{d}\mathbf{x}/\mathrm{d}t=\mathbf{v}_{\theta}$ if no interaction is considered, or with a modified velocity $\widetilde{\mathbf{v}}$ incorporating the interaction potential $\Phi$: $\widetilde{\mathbf{v}}(\mathbf{x}^{i},t)={\mathbf{v}}(\mathbf{x}^{i},t)-\frac{1}{N-1}\sum_{j\neq i}{k(\mathbf{x}^{i},\mathbf{x}^{j})}\nabla_{\mathbf{x}^{i}}\Phi(\mathbf{x}^{i}-\mathbf{x}^{j})$. Additionally, a weight mapping function $\phi_{\theta}^{g}$ models particle weights via $\mathrm{d}\log w_{i}(t)/\mathrm{d}t=g_{\theta}(\mathbf{x}_{i}(t),t)$, starting from initial weights $w_{i}(0)=1/N$ where $N$ represents batch size. Mathematically, we use the empirical measure $\mu_{0}^{N}=\frac{1}{N}\sum_{i=1}^{N}\delta_{x_{i}}$ to approximate the true distribution $\mu_{0}$, hence the uniform weights. This convergence to the true distribution is guaranteed when $N\rightarrow\infty$. The mass matching loss $\mathcal{L}_{\text{Mass}}$ is composed of two terms. The first term is defined as the local mass matching loss $\mathcal{L}_{\text{Local Mass}}=\sum_{k=1}^{T-1}M_{k}$, where $M_{k}$ quantifies the error between predicted weights $w_{i}(t_{k})$ and target weights based on the cardinality of mapped points. As detailed in Section˜A.4, the target weights, derived from the number of closest real data points, encourages the growth network to assign higher weights to particles moving into denser state space regions, thus provides fine-grained guidance on the growth network. Besides, the second term is defined as the global mass matching loss $\mathcal{L}_{\text{Global Mass}}=\sum_{k=1}^{T-1}G_{k}$, where $G_{k}$ is used to align the change of weights in total. An optimal transport loss $\mathcal{L}_{\text{OT}}=\sum_{k=1}^{T-1}\mathcal{W}_{2}(\hat{\mathbf{w}}^{k},\mathbf{w}(t_{k}))$ further aligns the predicted and observed distributions. Details can be found in Section˜A.4.
@@
CytoBridge aims to train four neural networks to model cell dynamics and interactions. To stabilize the training procedure, we leverage a two-phase training strategy. In the pre-training phase, we seek to provide a suitable initialization of these four neural networks. We first initialize $g_{\theta}$ and $\mathbf{v}_{\theta}$ without the interaction term to provide an approximated matching. Then we train ${\Phi}_{\theta}$ with fixed $g_{\theta}$, leading to refined dynamics. The score network $s_{\theta}$ is trained based on conditional flow matching. In the training phase, these initialized networks are further refined by minimizing the proposed total loss. We summarize the training procedure in Algorithm˜2 and Appendix˜A. We conduct ablation studies on different components of our training procedure in Section˜B.7. The selection of loss weighting is discussed in Section˜C.2.
@@
1: Datasets $A_{0},\ldots,A_{T-1}$, batch size $N$, ODE iteration $n_{\text{ode}}$, log density iteration $n_{\text{log-density}}$
@@
2: Trained neural ODE $\mathbf{v}_{\theta}$, growth function $g_{\theta}$, score network $s_{\theta}$ and interaction ${\Phi}_{\theta}$. 3:
4: Pre-Training Phase:
5: for $i=1$ to $n_{\text{ode}}$ do $\triangleright$ 1. Initialize growth $g_{\theta}$ and velocity $v_{\theta}$
6:   for $t=0$ to T2T-2 do
7:    $\hat{A}_{t+1}\leftarrow\phi_{\theta}^{\mathbf{v}}(\hat{A}_{t},t+1)$, $w(\hat{A}_{t+1})\leftarrow\phi_{\theta}^{g}(w(\hat{A}_{t}),t+1)$
@@
8:    $\mathcal{L}_{\text{Recons}}\leftarrow\lambda_{m}M_{t}+\lambda_{d}\mathcal{W}_{2}(\hat{\mathbf{w}}^{t},\mathbf{w}(t))$; update $\mathbf{v}_{\theta},g_{\theta}$ w.r.t. $\mathcal{L}_{\text{Recons}}$
9: for $i=1$ to $n_{\text{ode}}$ do $\triangleright$ 2. Initialize interaction potential $\Phi_{\theta}$
10:   for $t=0$ to T2T-2 do
11:    $\hat{A}_{t+1}\leftarrow\phi_{\theta}^{\widetilde{\mathbf{v}}}(\hat{A}_{t},t+1)$, $w(\hat{A}_{t+1})\leftarrow\phi_{\theta}^{g}(w(\hat{A}_{t}),t+1)$
12:    $\mathcal{L}_{\text{Recons}}\leftarrow\mathcal{W}_{2}(\hat{\mathbf{w}}^{t},\mathbf{w}(t))$; update $\mathbf{v}_{\theta},\Phi_{\theta}$ w.r.t. $\mathcal{L}_{\text{Recons}}$
13: for $t=0$ to T2T-2 do $\hat{A}_{t+1}\leftarrow\phi_{\theta}^{\mathbf{v}}(\hat{A}_{t},t+1)$
                                  $\triangleright$ Generate datasets $\hat{A}_{t}$.
14: for $i=1$ to $n_{\text{log-density}}$ do $\triangleright$ 4. Initialize the score network $s_{\theta}$
15:   $(\mathbf{x}_{0},\mathbf{x}_{1})\sim q(\mathbf{x}_{0},\mathbf{x}_{1})$, $t\sim\mathcal{U}(0,1)$, $\textbf{x}\sim p(\mathbf{x},t\mid\mathbf{x}_{0},\mathbf{x}_{1})$ using $\hat{A}_{0},\ldots,\hat{A}_{T-1}$
@@
16:   $\mathcal{L}_{\text{score}}\leftarrow\|\lambda_{s}\nabla_{\mathbf{x}}s_{\theta}(\mathbf{x},t)+\bm{\epsilon}_{1}\|_{2}^{2}$; update $s_{\theta}$ w.r.t. $\mathcal{L}_{\text{score}}$
17:
18: Training Phase:
19: Estimate initial distribution $p_{0}(\mathbf{x})$ from $A_{0}$ using Gaussian Mixture Model (GMM).
20: for $i=1$ to $n_{\text{ode}}$ do
21:   for $t=0$ to T2T-2 do
22:    $\hat{A}_{t+1}\leftarrow\phi_{\theta}^{\widetilde{\mathbf{v}}}(\hat{A}_{t},t+1)$, $w(\hat{A}_{t+1})\leftarrow\phi_{\theta}^{g}(w(\hat{A}_{t}),t+1)$
23:    $\mathcal{L}_{\text{Energy}}\leftarrow\mathbb{E}_{\mathbf{x}_{\tau}\sim\rho_{\tau}}\int_{\tau=t}^{\tau=t+1}\left[\frac{1}{2}\left\|{\mathbf{v}}_{\theta}\right\|_{2}^{2}+\frac{1}{2}\left\|\nabla_{\mathbf{x}}s_{\theta}\right\|_{2}^{2}\\ +\|{\mathbf{v}}_{\theta}\|_{2}\|s_{\theta}\|_{2}+\alpha\Psi\left(g_{\theta}\right)\right]w_{\theta}(\tau)\mathrm{d}\tau$
24:    $\mathcal{L}_{\text{Recons}}\leftarrow\lambda_{m}(M_{t}+G_{t})+\lambda_{d}\mathcal{W}_{2}(\hat{\mathbf{w}}^{t},\mathbf{w}(t))$
25:    $\mathcal{L}_{\text{FP}}\leftarrow\left\|\partial_{\tau}\rho_{\theta}(\mathbf{x},\tau)+\nabla_{\mathbf{x}}\cdot\left(\rho_{\theta}(\mathbf{x},\tau)\widetilde{\mathbf{v}}_{\theta}(\mathbf{x},\tau)\right)-g_{\theta}(\tau)\rho_{\theta}(\mathbf{x},\tau)\right\|+\lambda_{w}\left\|\rho_{\theta}(\mathbf{x},0)-p_{0}(\mathbf{x})\right\|$
26:    $\mathcal{L}_{\text{Total}}\leftarrow\mathcal{L}_{\text{Energy}}+\lambda_{r}\mathcal{L}_{\text{Recons}}+\mathcal{L}_{\text{FP}}$; update $\mathbf{v}_{\theta},g_{\theta},s_{\theta},\Phi_{\theta}$ w.r.t. $\mathcal{L}_{\text{Total}}$
6 Results
Next, we evaluate CytoBridge’s ability to simultaneously learn cell dynamics and cell-cell interactions. In computations below, we take $\Psi(g(\mathbf{x},t))=g^{2}(\mathbf{x},t)$ and $\sigma(t)$ is constant. We also assume the interaction term is dependent on the distance between cells in gene expression space and we use radial basis functions (RBFs) to approximate it (Section˜A.3).
Synthetic Gene Regulatory Network
In order to examine CytoBridge’s capabilities of learning cell dynamics as well as their underlying interactions simultaneously, we conducted experiments on the three-gene simulation model following (Zhang et al., 2025a). The dynamics of the original three-gene model are governed by stochastic ordinary differential equations, incorporating self-activation, mutual inhibition, and external activation (Section˜B.1), as shown in Fig.˜2 (a). Additionally, we incorporated interactions into the simulation process. We consider the following types of interactions: (1) attractive interactions. (2) Lennard-Jones-like potential (both attractive and repulsive) (3) no interactions. We aim to test whether CytoBridge can recover interaction in each case. For attractive interactions, the cells with similar gene expressions tend to converge toward similar levels. The interaction potential $\Phi$ is defined as: $\Phi(\mathbf{x}-\mathbf{y})=\left\|\mathbf{x}-\mathbf{y}\right\|^{2}$ As shown in Fig.˜2(b), the dynamics of the three-gene model exhibit a quiescent area as well as an area with notable transition and increasing cell numbers. Moreover, the attractive potential results in the reduction in variance of the observed data at different time points. We compared CytoBridge with other methods across all time points using the Wasserstein distance ($\mathcal{W}_{1}$) and the Total Mass Variation (TMV) metric, defined in Appendix˜C. We summarized the results in Table˜1. It is shown that CytoBridge achieves the best performance in both distribution matching and mass matching. Balanced Schrödinger bridge (e.g., (Tong et al., 2024b)), which neglects both growth and interaction terms, results in false transition and variance patterns (Fig.˜2(c)). DeepRUOT (Zhang et al., 2025a), which is an unbalanced Schrödinger Bridge solver, exhibits correct transition patterns but fails to capture the reduction in variance as it neglects the cell-cell interactions (Fig.˜2(d)). By leveraging the UMFSB framework, as shown in Fig.˜2(e), CytoBridge correctly models both the transition patterns and the reduction in variance as the correct interaction potential (Fig.˜2(f)) and growth rate (Fig.˜2(g)) can be directly learned by CytoBridge. Also, CytoBridge is capable of constructing the underlying Waddington landscape (Fig.˜2(h)). The low-lying regions on this landscape correspond to areas of high probability density, representing stable states. Other than the attractive interaction potential, we also incorporated the Lennard-Jones-like potential, and no interaction case. Results can be found in Section˜B.1, Figs.˜4 and 5 and we find CytoBridge can correctly identify both the LJ potential and no interaction in these cases. Overall, both quantitative results and qualitative analysis indicate the necessity of incorporating both growth and interaction terms.
Figure 2: (a) Illustration of the synthetic gene regulatory dynamics. (b) The ground truth cellular dynamics project on $(X_{1},X_{2})$. The red lines indicate the ground truth trajectories of cells in (b), or inferred trajectories of cells in (c) to (e). (c) The dynamics learned by balanced Schrödinger bridge SF2M (Tong et al., 2024b). (d) The dynamics learned by DeepRUOT. (e) The dynamics learned by CytoBridge. (f) The learned interaction potential. (g) The growth rates inferred by CytoBridge. (h) The constructed landscape at $t=4$. The z-axis represents the density of cells.
$t=1$ $t=2$ $t=3$ $t=4$